Solve the equation for x. 5x - 3 = 4 - 3x
Found 3 solutions by Tobiasz, Alan3354, richwmiller:
Answer by Tobiasz(54)

5x - 3 = 4 - 3x
7x-3=4
-3=4-7x
1= 7x

x=0.14

Answer by Alan3354(69443)

5x - 3 = 4 - 3x
8x = 7
x = 7/8

Answer by richwmiller(17219)

5x - 3 = 4 - 3x
5x+3x=7
8x=7
x=7/8